Science Report – December 26th

26 December 2018
Crew 201 – Mex-1 Crew
Submitted by Federico Martínez, Crew Astronomer/Scientist

1. The Multidimensional Fatigue Symptom Inventory: The crew is responding to a questionnaire every day, and creating a database for its future analysis.

2. Crew Wellness Experiment: Between 7 am and 8 am, every member of the crew has been measuring its weight and pressure, and taking notes of water consumption when the day is over. All of the crew members have lost weight, between 500gr and 1.5 kgs.

3. Very Small Aperture Terminal (VSAT) Pointing: The spaceship that carries on the Satellite system has been lost; a new mission has been launched to deliver a new satellite system, awaiting confirmation to receive the new system.

4. 3D Printing in space exploration: 3D printer it’s been used to make tools and the prototyping of a Rover. The cold weather makes the correct functionality of the printer in the RAM difficult, so we decided to move to the first floor of the Hab.

5. Engaging space to the people: Material has been recorded on each one of the MDRS modules. To complete this project, we just need to record the personal videos of each member.

6. Validation of electronics architecture and communication protocols for an exploration rover: Waiting for the transmission module to arrive at the station, so we can test next Friday with the mechanical prototype.

7. The behavior of Artificial Vision algorithms for Autonomous Navigation: The necessary images to train the algorithm have been taken, but we are having trouble with software settings.

8. Prototype and mechanical testing of Exploration rover: The rover has been 90% assembled and we are printing the last pieces so we can fully test on Friday around the Habitat.

9. Martian Soil Analysis for usage on Greenhab: Categorization of soil based on colors and grain size, tomorrow will proceed with CO, Ph, and salinity measurements.

Science Report – December 19th

MS,

This is our Science Report until today
Thanks!

Science Report 19 Dec 2018
Crew 201 – Mex-1 Crew
Submitted by Crew Astronomer/Scientist Federico Martínez

1. The Multidimensional Fatigue Symptom Inventory: The crew is responding a questionary every day, and creating a database for its future analysis.

2. Crew Wellness Experiment: Between 7 am and 8 am, every member of the crew has been measure its weight and pressure, and taking notes of water consumption when the day is over.

3. Very Small Aperture Terminal (VSAT) Pointing: The spaceship that carries on the Satellite system is going to arrive tomorrow at the station.

4. 3D Printing in space exploration: 3D printer fully assemble and tested, it arrived to the station with malfunction, two days of a assemble and repair, now it works correctly.

5. Engaging space to the people: Work in progress.

6. Validation of electronics architecture and communication protocols for an exploration rover: Currently, making electronics and communication tests, working on the communication between RF modules and computer.

7. Behavior of Artificial Vision algorithms for Autonomous Navigation: Having trouble with the setup of the camera recognition software, the team went to an EVA to collect picture samples to train the algorithm.

8. Prototype and mechanical testing of Exploration rover: Work in progress.

9. Martian Soil Analysis for usage on Greenhab: First day of samples recollection, testing will begin on Dec 20.

Science Report – May 04th

Here is the Science Report for 4 May 2018
Crew 193 – PHEnOM Gold Crew
Submitted by XO/Science Officer Doug Campbell

1. Emergency EVAC EVA: No further emergency evacuation locations were scouted.

2. Sunspot and Solar Flare Monitoring: One solar flare was spotted today on the sun’s southwestern side.

3. Shortwave Texting on EVA’s: Beartooth devices are still not working and were not taken on EVA today.

4. GPS Route Measurements: Both EVAs today took the GPS with them to track their routes for their PHEnOM colleague.

5. Waterless dish cleaning: Bacteria growth and monitoring came to an end today. The study showed promising results regarding the use of UV and a cloth to clean dishes on mars without the use of water.

6. The MAG (Maximum Absorbency Garment) utilization study: The study has been scrapped. Results will be from blue crew only.

7. Spacesuit Helmet Fogging: Study has been discontinued after 7 days of data collection.

8. Crew Comfort, Health and Safety on EVA: Final measurements were taken for all crew members during EVA today. Analysis of the data will occur after the mission is completed.

Thanks,
Doug
Crew 193 XO and Science Officer

Science Report – 03 May 2018

1. Emergency EVAC EVA: No EVAs occurred today due to visit from an elementary school class.

2. Sunspot and Solar Flare Monitoring: No pictures were taken with the camera today, because we hosted a group of school kids on a tour of the entire campus.

3. Shortwave Texting on EVA’s: No EVAs occurred today due to visit from an elementary school class.

4. GPS Route Measurements: No EVAs occurred today due to visit from an elementary school class.

5. Waterless dish cleaning: Bacteria growth and monitoring will continue over the last two days of the mission. The study is continuing to show promising results regarding the use of UV and a cloth to clean dishes on mars without the use of water.

6. The MAG (Maximum Absorbency Garment) utilization study: The study has been scrapped. Results will be from blue crew only.

7. Spacesuit Helmet Fogging: Study has been discontinued after 7 days of data collection.

8. Crew Comfort, Health and Safety on EVA: No EVAs occurred today due to visit from an elementary school class.

Science Report May 2nd

Hello mission control,

Here is the Science Report for 2 May 2018

Crew 193 – PHEnOM Gold Crew

Submitted by XO/Science Officer Doug Campbell

1. Emergency EVAC EVA: No EVAs occurred today due to weather.

2. Sunspot and Solar Flare Monitoring: a very small prominence was found on the sun today.

3. Shortwave Texting on EVA’s: No EVAs occurred today due to weather

4. GPS Route Measurements: No EVAs occurred today due to weather

5. Waterless dish cleaning: Bacteria growth and monitoring will continue over the remaining days of the mission. The study is starting to show promising results regarding the use of UV and a cloth to clean dishes on mars without the use of water. Using dirt to clean the dishes seems to have some rapid growing bacteria!

6. The MAG (Maximum Absorbency Garment) utilization study: The study has been scrapped. Results will be from blue crew only.

7. Spacesuit Helmet Fogging: Study has been discontinued after 7 days of data collection.

8. Crew Comfort, Health and Safety on EVA: No EVAs occurred today due to weather

.

Thanks,

Doug

Crew 193 XO and Science Officer

Science Report – May 1st

Crew 193 – PHEnOM Gold Crew

Submitted by XO/Science Officer Doug Campbell

1. Emergency EVAC EVA: Two EVAs took place today; however, no new sites were noted as useful for emergency evacuation shelters.

2. Sunspot and Solar Flare Monitoring: One small, shallow prominence was photographed on the Sun. No other features were apparent.

3. Shortwave Texting on EVA’s: Beartooth devices are still not working. Troubleshooting through contact with the manufacturer is occurring.

4. GPS Route Measurements: Both EVAs took the device with them to map their route. Measurements will be used by a PHEnOM colleague to map the routes taken and determine if the easiest routes were used.

5. Waterless dish cleaning: Bacteria growth and monitoring will continue over the remaining days of the mission. The study is starting to show promising results regarding the use of UV and a cloth to clean dishes on mars without the use of water. Using dirt to clean the dishes seems to have some rapid growing bacteria!

6. The MAG (Maximum Absorbency Garment) utilization study: The study has been scrapped. Results will be from blue crew only.

7. Spacesuit Helmet Fogging: Study has been discontinued after 7 days of data collection. No fogging was ever reported. One helmet fogged today on EVA during a brief rainfall. Once rain stopped, it returned to un-fogged almost immediately.

8. Crew Comfort, Health and Safety on EVA: Measurements of the crew’s weight, blood pressure, SpO2 and pulse were taken before and after the EVA. Maximum heart rate was tracked by a mobile monitor during the EVA when the astronaut felt that they were working hard (normally climbing a large hill. Analysis will be done as the mission continues.

Science Report – April 30th

Crew 193 – PHEnOM Gold Crew

Submitted by XO/Science Officer Doug Campbell

1. Emergency EVAC EVA: Two EVAs took place today; however, no new sites were noted as useful for emergency evacuation shelters. The second EVA found a location in Candor Chasma that had been previously scouted by the other EVA team last week.

2. Sunspot and Solar Flare Monitoring: The crew astronomer saw one filament on the sun. No other features were apparent.

3. Shortwave Texting on EVA’s: Beartooth devices are still not working. Troubleshooting through contact with the manufacturer is occurring.

4. GPS Route Measurements: Both EVAs took the device with them to map their route. Measurements will be used by a PHEnOM colleague to map the routes taken and determine if the easiest routes were used.

5. Waterless dish cleaning: Bacteria growth and monitoring will continue over the remaining days of the mission. The study is starting to show promising results regarding the use of UV and a cloth to clean dishes on mars without the use of water.

6. The MAG (Maximum Absorbency Garment) utilization study: The study has been scrapped. Results will be from blue crew only.

7. Spacesuit Helmet Fogging: Study has been discontinued after 7 days of data collection. No fogging was ever reported.

8. Crew Comfort, Health and Safety on EVA: Measurements of the crew’s weight, blood pressure, SpO2 and pulse were taken before and after the EVA. Maximum heart rate was tracked by a mobile monitor during the EVA when the astronaut felt that they were working hard (normally climbing a large hill. Analysis will be done as the mission continues.

Science Report – April 29th

Science Report for 29 April 2018
Crew 193 – PHEnOM Gold Crew
Submitted by XO/Science Officer Doug Campbell

1. Emergency EVAC EVA: Two EVAs took place today; however, no sites were noted as useful for emergency evacuation shelters. If a crew was trapped north on Copernicus highway, there are several low sites where they could shelter out of the wind.

2. Sunspot and Solar Flare Monitoring: Our CGA continued his task of monitoring the sun for sunspot activity. He found two dark features on the sun and another possible solar flare.

3. Shortwave Texting on EVA’s: Beartooth devices are still not working. Troubleshooting through contact with the manufacturer is occurring.

4. GPS Route Measurements: The first EVA of the day took the device with them to map their route. Measurements will be used by a PHEnOM colleague to map the routes taken and determine if the easiest routes were used.

5. Waterless dish cleaning: Final samples of the project were taken from a towel wiped dirty dish that had been place under UV and a dirty dish that had been wiped with dirt and then placed under a UV light. Bacteria growth and monitoring will continue over the remaining days of the mission.

6. The MAG (Maximum Absorbency Garment) utilization study: No progress today. Possible that the continuation of this study from Crew 192 will be scrapped.

7. Spacesuit Helmet Fogging: No fogging was reported.

8. Crew Comfort, Health and Safety on EVA: Measurements of the crew’s weight, blood pressure, SpO2 and pulse were taken before and after the EVA. Maximum heart rate was tracked by a mobile monitor during the EVA when the astronaut felt that they were working hard (normally climbing a large hill. Analysis will be done as the mission continues.

Thanks,
Doug
Crew 193 XO and Science Officer

Science Report – April 28th

Crew 193 – PHEnOM Gold Crew

Submitted by XO/Science Officer Doug Campbell

1. Emergency EVAC EVA: No EVAs took place today.

2. Sunspot and Solar Flare Monitoring: Our CGA continued his task of monitoring the sun for sunspot activity. He found a solar prominence during his imaging session this morning.

3. Shortwave Texting on EVA’s: No EVAs took place today.

4. GPS Route Measurements: No EVAs took place today.

5. Waterless dish cleaning: Samples were taken from a dirty dish that had been place under UV and a dirty dish that had been wiped with dirt and then placed under a UV light. Bacteria are growing on the samples from the first test.

6. The MAG (Maximum Absorbency Garment) utilization study: No progress today. Possible that the continuation of this study from Crew 192 will be scrapped.

7. Spacesuit Helmet Fogging: No EVAs took place today.

8. Crew Comfort, Health and Safety on EVA: No EVAs took place today.

Science Report – April 27th

Science Report for 27 April 2018
Crew 193 – PHEnOM Gold Crew
Submitted by XO/Science Officer Doug Campbell

1. Emergency EVAC EVA: No locations were scouted during the EVA today. Emergency triage procedures were practiced so that a crew member might be safely transported to an evacuation location in the event of an incident.

2. Sunspot and Solar Flare Monitoring: Our CGA continued his task of monitoring the sun for sunspot activity. He found a solar flare in the direction of our planet and immediately alerted mission control to the danger.

3. Shortwave Texting on EVA’s: Hand signal communication was used during the emergency EVA practice today. Further practice is needed to achieve a smooth process.

4. GPS Route Measurements: No measurements today, as the EVA was at the hab location.

5. Waterless dish cleaning: Samples were taken from a dirty dish and a dirty dish that had been put under the UV light. There are minimal signs of bacterial growth from the first set of samples. It may take a couple more days to see signs of growth. Both sets of samples are now in the incubator in the hab.

6. The MAG (Maximum Absorbency Garment) utilization study: No progress today. Possible that the continuation of this study from Crew 192 will be scrapped.

7. Spacesuit Helmet Fogging: Neither crew experienced fogging of the masks today. Both crews used a defogging spray and non-scratch cloth on the masks prior to their EVAs today.

8. Crew Comfort, Health and Safety on EVA: Measurements of the crew’s weight, blood pressure, SpO2 and pulse were taken before and after the EVA. Analysis will be done as the mission continues.